OpenStudy (anonymous):

|\sqrt{2} - 1| When you remove the absolute bars from this, you get: |\sqrt{2} + 1|, right? |\sqrt{2} - 1| When you remove the absolute bars from this, you get: |\sqrt{2} + 1|, right? @Mathematics

OpenStudy (anonymous):

I meant: \sqrt{2} + 1

OpenStudy (anonymous):

no - but it can be sqrt(2) - 1 or -(sqrt(2) - 1)

OpenStudy (anonymous):

How?

OpenStudy (jamesj):

By definition |x| = x, if x > 0 0, if x = 0 -x, if x < 0

OpenStudy (jamesj):

Now \( \sqrt{2} - 1 > 0 \) hence \( | \sqrt(2) - 1 | = \sqrt{2} - 1 \).

OpenStudy (jamesj):

What perhaps you're thinking of is this: \[\frac{1}{\sqrt{2}-1} = \frac{1}{\sqrt{2}-1} . \frac{\sqrt{2}+1}{\sqrt{2}+1} = \frac{\sqrt{2}+1}{2 - 1} = \sqrt{2} + 1\]